OpenStudy (anonymous):

. Indicate which property is illustrated in Step 2. Step 1 3 + (1 + 7) + 0 = 3 + (7 + 1) + 0 Step 2 = (3 + 7) + (1 + 0) Step 3 = (3 + 7) + 1 Step 4 = 10 + 1 A. commutative property of multiplication B. associative property of addition C. identity property of addition D. associative property of multiplication

OpenStudy (anonymous):

These are all simple.. Commutative multiplication you need to find a 'x' in order for it to be multiplication. Associative property of addition is where you add 2 sums and the answer is the same if you flip the equation. (5 + 2 = 7 = 2 + 5). Identity property of addition is when the equation has a zero in it and the answer is the same as the sum in the equation. ( 6 + 0 = 6) And associative property of multiplication is where three or more numbers are multiplied. Do you understand those??

OpenStudy (anonymous):

yeah that why I got b

OpenStudy (anonymous):

Ok. B is not right. We are looking at step 2, (3 + 7) + (1 + 0). associative property of addition is where you add 2 sums ( 5 + 2 = 7) and the answer comes out the same even when you flip it. ( 2 + 5 = 7).

OpenStudy (anonymous):

well it cant be identity can it

OpenStudy (anonymous):

You can't just assume it's an answer. You have to do the work to solve the answer. Let me look at the problem again.

OpenStudy (anonymous):

(3 + 7) can you add that and get the same number if you flipped it?